West Coast Tours runs boat tours along the west coast of British Columbia. On March 5, 2017, it purchased, with cash, a cruising boat for $894,000, having a useful life of 10 years or 16,725 hours with a residual value of $225,000. The company’s year-end is December 31.

Required:
Calculate depreciation expense for the fiscal years 2017, 2018, and 2019 by completing a schedule. (Note: Depreciation is calculated to the nearest month. Assume actual hours of service were: 2017, 830; 2018, 1,890; 2019, 1,645.)

Depreciation Method
Year Straight-Line Double-Declining Balance Units-of-Production
2017
2018
2019
Answer #1
Year Straight Double Units of
line declining production
2017 55,750 149,000 33200
2018 66,900 149,000 75600
2019 66,900 119,200 65800
Straight line = (894000-225000)/10
66900
for first year dep will be for march to december 66900/12*10= 55750
Double Declining method
1/10*2= 0.2 or 20%
2017 894000*20%= 178800 *12/10= 149000
2018 (894,000-149,000)*20%= 149000
2019 (894000-149000*2)*20%= 119200
units of production
depreciation rate = ( 894000-225000)/16725
40